# Teleoscope: Exploring Themes in Large Document Sets By Example

### v1

@@ 4. Design Studies
This section outlines our design study processes. We evaluated Teleoscope through (1) informal piloting and internal analysis using low-cost evaluation methods; (2) a multi-week study and focus group with qualitative researchers (N=5) including a post-hoc expert review of our interface and study data with a visualization group; (3) a multi-month field deployment with a qualitative research group; and (4) an on-going public release. We used data from Reddit as archived by PushShift (Baumgartner et al., 2020) up to their latest published data in January 2023. Except for initial in-lab piloting, all research was conducted under approval of our institution’s research ethics board and all participants signed initial consent forms and were reminded of their right to halt participation in the studies if desired.
@@
We focused on one subreddit, the r/AmItheAsshole advice forum since we were ourselves interested in data on social norms; our field deployment collaborators also wanted access to r/nursing. r/AmItheAsshole has roughly 650K documents and r/nursing has roughly 100K documents. Data included only posts, not comments. The last post date was February 2023, since Reddit significantly restricted their data API access in response to ChatGPT.
4.1. Design Study 1: Initial piloting and internal analysis
@@
Teleoscope was initially developed for our own use as a research tool for searching through large corpora of text and went through a series of very quick design iterations with a single designer. As we developed it into a more robust and large-scale system, we began to incorporate more human-centred design techniques into our design process as our research and development team grew. Since this was early stage design, the interface design changed significantly as we determined design requirements. We are including only light details on our informal methods so as to faithfully report on our process. During this year-long phase, we also ran a series of tests on different NLP approaches. These were incorporated into our system during our informal user testing.
@@ 4.1.3. Results
Recognition over recall. By moving commands and system state out of menus/collapsible dashboard modules, we opted for a design with minimum display of information. This is contrary to an approach that attempts to display the full system state through statistics or summary graphics. We deemed this unnecessary for our core interaction goal; opting instead for an Overview/Details-on-Demand design pattern where the process was visualized rather than the full system state.
@@
4.1.4. Design Study 1: Conclusion
During this study, we quickly worked through interaction and visualization approaches that are common in topic modelling. After informal user feedback and using the system ourselves, we started to develop interaction methods that focused on directly manipulating data provenance.
4.2. Design Study 2: Multi-week user study and Focus Group
@@
  • E2.RQ1. In what ways did participants understand and use features such as collaboration, search and Teleoscope ranking?
  • E2.RQ2. How did participants incorporate Teleoscope into their understanding of qualitative research processes?
@@ 4.2.4. Diary Results
Teleoscoping differs from keyword searches. Participants found that the Teleoscope ranking system differed from a normal keyword search. For many participants, it took some time to (1) differentiate results of a keyword search from results of a Teleoscope rank operation; and (2) differentiate valid results that did not meet their expectations due to the documents that existed in the corpus from invalid results due to bugs or mental model inconsistencies. There were negative transfer effects from being used to keyword searches which took multiple sessions to unlearn.
For example, P1 searched for “palliative” and was “…surprised by how many posts were about animals at end of life, which does not fit our defined research topic.” P1 then wished “…there was a feature that would take everything I had already put within one group and give me ‘more like this’”, which was exactly what the Teleoscope rank operation was designed to do. Multiple participants corroborated this sentiment in their diaries (P5, P7, P8). This indicated a problem in the participants’ mental model, likely due to (1) how we were representing the Teleoscope operation on the workspace; (2) our training, documentation and support materials; and (3) not enough time to learn the tool.
@@
Today I also spent time trying to go down the rabbit hole of different searches to try to gain a true appreciation for how this machine learning approach to data collection differs from just keyword searching within the Reddit search. This was really evident to me when I found a post where the OP had a palliative/terminal illness, and I wanted to find others where this was the case. I made a new folder for this category, then used the Teleoscope feature, and immediately found one other post where the OP has cancer and was asking a friend to not mourn their death. It would be extremely difficult to keyword search for this type of topic, but it’s a very interesting and important area to capture (OPs with terminal illnesses). This was a great exploration! (P1)
This indicated that it was possible to learn the difference between keyword search and Teleoscope rank, but that the learning curve was steep enough to require multiple hours of usage and documentation review.
@@
They further suggested adding a document quick viewer to aid in skimming. They then organized documents into groups, relabelling and changing the groups as they developed their understanding of the corpus and the tool. Then, they switched between using the teleoscope operation and the group feature to find relevant documents:
@@
After adding about half a dozed documents to the group ‘gay stuff’, I noticed that many of the documents are about gay panic. That is, the fear of being (wrongly or correctly) as gay, the dislike of anything perceived as gay, and an aversion from being around gay people. I changed the group name to ‘gay panic‘ to reflect this…once I had about 13–14 results, I opened the Teleoscopes window for the group. Looking at the first two pages of results, none of them that weren’t already in the group seemed very relevant, mostly judging by their titles and occasionally by the first couple of sentences in the document. I refined the search…[with a] couple of documents that I thought particularly demonstrated gay panic. (P7)
@@ 4.2.5. System Log Results
During the user study, the system maintained a log of user actions as they interacted with the system. Across the 5 participants, 7 sessions with the teleoscope operations were tracked (one participant created 3 separate sessions for themselves). The mean number of actions tracked per session was approximately 310 (median = 286, minimum = 97, maximum = 684). Actions included such things as session initialization, creation/movement/deletion of windows, keyword searches, instantiation of teleoscope ranking and results, in short, any conditions where the state of the user workspace was altered.
The actions were logged and then visualized to better understand user interactions. Generally, users made use of an iterative process to find documents of interest, alternating between putting documents into groups and instantiating new teleoscopes to find new documents relevant to their queries and then sorting them into their groups. The number of teleoscopes created across the study were relatively small however (between 2 and 4).
@@ 4.2.6. Focus Group Results
Mental model need: teleoscope state needs to be inspectable. In our tested design, teleoscope state was coupled directly to a single group. When the group was updated, the teleoscope rank changed. Individual documents within the group could be selected to weight the teleoscope search vector closer to that document. This confused participants about the state of the teleoscope rank. We decided that the teleoscope inputs and outputs needed a more explicit visual representation and decoupled from a single group, allowing for multiple input sources.
@@
Mental model need: Clarify the Teleoscope exploration metaphors onscreen. For example, participants wondered whether they were being “dropped off in the landscape” of documents and “going down different paths” (P7). If so, they wanted a record of the paths and some way to compare paths directly onscreen. Participants agreed that “seeing it” and “understanding how close documents are to each other in space [is important].” (P1)
@@
Conceptual need: Confidence in path saturation. Participants agreed that they did not need to show total path exhaustion, rather, they needed a sense of saturation. Their imagination of the use case for Teleoscope was for data collection, which did not mean finding every piece of relevant data but instead finding enough representative samples of data.
Conceptual need: Explaining the methodology to reviewers. Participants were concerned about how to explain the Teleoscope process to reviewers at a high level, but were not concerned with the details of the Teleoscope process. As long as they had a clear metric that they could point reviewers to (e.g., a paper that describes the metric), they did not particularly care about which metric to use. This was a defensive publication strategy: keyword searches are reported on directly to show paper saturation but there seems to be an understanding among current reviewers that the immense size of possible data sets meant that any reasonable metric could be argued for and used.
Conceptual need: differentiating data collection from analysis. Participants were unsure whether using Teleoscope constituted a violation of rigour such that it mixed data collection and analysis research phases. Particularly if text-level analysis was to be supported, they felt that it might not be appropriate to allow for both in the interface at once. We found that participants used Teleoscope as we intended: they explored data in a manner that extended beyond keywords searches. For example, participants reported the following:
@@ 4.2.7. Post-Hoc Analysis and Recommendations from Visualization Group
4.2.8. Design Study 2: Conclusion
@@
  • E2.RQ1. Participants understood the overall concept of the Teleoscope ranking operation and workspace as a whole, likening it to walking down different paths in an unknown region. Even though this metaphor came from the participants, the operation inputs and system state were unclear; therefore we decided that the path metaphor should be visualized as directly as possible. Participants did not use any collaboration features of Teleoscope, which may be due to a lack of impetus, our study environment, or because of our design. In the next study, we addressed these possibilities to ensure collaboration features were used.
  • E2.RQ2. Participants interpreted Teleoscope primarily as a data collection tool. Further, they were unsure about whether it infringes on analysis and had concerns about academic rigour if it does.
4.3. Design Study 3: Multi-Month Field Deployment
After completing a redesign after Study 2, we were interested in a long-term focused field deployment with qualitative research groups who could bring their own research needs to us. We also wanted to move Teleoscope out of a simulated environment and into a production environment where real-world motivations and difficulties would be encountered.
Study 2 had many study contrivances, such as a research topic that none of the participants were using for their own research. We felt that it was important to see how Teleoscope would perform when participants were subjected to all the benefits, consequences and costs of real research. We also committed to ongoing feature development (and bug fixes) using a quasi co-design methodology.
@@ 4.3. Design Study 3: Multi-Month Field Deployment
  • E3.RQ1. When put in a real-world environment, how did researchers incorporate Teleoscope into their own research practice? Which features, processes, and workflows within Teleoscope were commonly used and in what way?
  • E3.RQ2. Using Teleoscope, to what extent were researchers able to feel confident that they were able to retrieve data using criteria that are important to qualitative researchers, i.e., richness, saturation, etc.?
@@ 4.3.1. Participants
Three PIs and research groups were recruited; only one research group was able to commit to long-term use of Teleoscope. The final team was comprised the PI from Study 2 and a research team of two graduate RAs that were recruited specifically to use Teleoscope. Participants were not reimbursed for their time by us since they were being supported extensively by our design team for their own research project; our understanding is that RAs were compensated as normal by the PI.
@@ 4.3.3. Results
Figure 9. An example of a workflow from Design Study 3 our long-term deployment. Pictures are actual screenshots of research artifacts from our participant research team’s data-gathering phase. Participants worked both individually and collaboratively on the Teleoscope interface, and collaboratively on Google Docs, Zoom, and in-person. Due to the existing qualitative research culture in their research group, keyword searches were the focus of their data collection approach. ML features were used to discover new keywords, find thematically similar documents that did not have specific keywords, and to saturate groups with relevant documents.
The research team used Teleoscope for data collection for their research project on nurses and structural inequality as articulated in Reddit’s r/nursing forum, which is where working nurses post about their day-to-day problems. Of their own initiative and in alignment with their own qualitative methodological approach, the team’s main conceptualization of data collection was via an external google document that contained a set of keywords to be used in searches (see Figure 9). After exploring the data using a variety of features in the Teleoscope workspace (described below), they would update their keyword list with successful and unsuccessful searches. Here are the ways in which the research team used the Teleoscope workspace features:
@@
Teleoscope helped with search saturation. The researchers reported (1) making groups from documents from their keyword searches; (2) piping the groups into the Teleoscope ranking operation as controls; (3) determining which documents they had not yet read and (4) adding those documents to the groups. This helped to see the parts of the document space that they had not yet captured with a keyword search.
@@
Putting each one of our single groups like indigenous, vulnerable disabilities, [an RA] put them into the teleoscope and then basically went through to see like which ones we hadn’t read…We were just trying to expand our data set and be exhaustive. (P1)
@@
It was actually really helpful to start with keyword searches. And to be able to build out this groups structure, and then Teleoscope from there. Whereas maybe if we’d had a more like drilled in topic, we could have just gone from there.(P1)
@@
Projections have potential to allay methodological concerns. In the focus group, it was brought up as a concern that using Teleoscope might be too close to analysis and confuse methodological rigour (in rigorous qualitative research, data collection and analysis stages are kept distinct so as not to predetermine results). However, the projection operation, which was added after the focus group, seemed to have potential to populate the interface with unexpected results:
@@
The speed and ease of use helped with the sense of completion as well:
@@
4.3.4. Design Study 3: Conclusion
@@
  • E3.RQ1. Researchers incorporated Teleoscope into their research practice as a data collection tool, working between an external Google doc for keywords and the interface itself. Teleoscope was used to explore parts of the document space where keywords would not be easy or obvious to find. The most common features used were the keyword search, document reading and grouping. Teleoscopes were used after grouping. Projections were used after Teleoscoping and grouping.
@@
  • E3.RQ2. Teleoscope helped to provide confidence that a corpus was being more fully and rigorously explored by providing both ranked and randomized example documents.
4.4. Design Study 4: On-going Public Release
We are hosting an ongoing public beta release of Teleoscope (publicly deployed at [anonymized for review after consultation with CHI chairs]). This is not a formal user study; instead it is an ongoing test and demonstration of our system robustness in terms of performance, security, and availability. By committing to a live release on a cloud platform, we were forced to develop the following security and availability measures:
  • (Nearly) one-click deployment. Teleoscope can be deployed on a new AWS virtual machine using Ansible playbooks in nearly one click. This was developed after a sprinkler accident destroyed our original non-public servers and motivated our move to the cloud. We needed to re-deploy Teleoscope often enough to spend time developing an automatic deployment system.
  • Robust backup system. We expected a catastrophic security breach at some point and developed a backup system. When our database was indeed hacked and erased, we redoubled our backup system to two small-scale hourly offsite backups as well as a daily backup.
  • User roles/API limiting/Reverse proxy/SSL/TLS. Earlier this year, Reddit restricted usage of the data API, which interfered with our data collection strategy. It also introduced the threat of large-scale data scraping from Teleoscope. As such, we restricted Teleoscope to registered users (open to anyone to register), creating a robust internal/external user role scheme for MongoDB and RabbitMQ, put in place data security measures such as limiting our API throughput, and set up a SSL/TLS reverse proxy to encrypt messages between client and server.
Teleoscope remains online as of the writing of this paper and continues to gain users.
5. Discussion, Limitations and Future Work
In this section, we discuss the results of our studies with regards to our design goals, research questions, and provide directions for future work.

### v2

@@ 4. Deployment Case Studies
This section outlines our evaluation processes. We evaluated Teleoscope through (1) informal piloting and internal analysis using low-cost evaluation methods; (2) a multi-week study and focus group with qualitative researchers (N=5) including a post-hoc expert review of our interface and study data with a visualization group; (3) a multi-month field deployment with a qualitative research group; and an on-going public release as we continue to develop the tool. We used data from Reddit as archived by PushShift (Baumgartner et al., 2020) up to their latest published data in January 2023. Research was conducted under approval of our institution’s research ethics board; participants signed consent forms and were reminded of their right to halt participation in the studies if desired.
@@
We focused on one subreddit, the r/AmItheAsshole advice forum since we were ourselves interested in data on social norms; our field deployment collaborators also wanted access to r/nursing and eventually their own arbitrary data uploads. r/AmItheAsshole has roughly 650K documents and r/nursing has roughly 100K documents. Data included only posts, not comments. The last post date was February 2023, since Reddit significantly restricted their data API access in response to ChatGPT.
4.1. Case Study 1: Piloting
@@
In the first year, Teleoscope first went through a series of very quick design iterations within our team. As we developed it into a more robust and large-scale system, we began to incorporate more human-centred design techniques into our design process as our research and development team grew. We are including only light details on our informal methods so as to faithfully report on our process. During this year-long phase, we also ran a series of tests on different NLP approaches. These were incorporated into our system during our informal user testing.
@@ 4.1.3. Results
Recognition over recall. By moving commands and system state out of menus/collapsible dashboard modules, we opted for a design with minimum display of information, contrary to our prior approach that displayed summary statistics of keyword distributions. We deemed this unnecessary for our core interaction goal; opting instead for an Overview/Details-on-Demand design pattern where the process was visualized rather than the full system state.
@@
4.1.4. Case Study 1: Conclusion
During this study, we quickly attempted and discarded visualization approaches that are common in topic modelling, including displaying and interacting most directly with keywords and topic labels. Instead, we moved towards a direct manipulation approach, where documents and outputs were visualized as arrangable windows on a workspace.
4.2. Case Study 2: User study and Focus Group
@@
  • CS2.RQ1. In what ways did participants understand and use features such as collaboration, search and ranking?
  • CS2.RQ2. How did participants incorporate Teleoscope into their understanding of qualitative research processes?
@@ 4.2.4. Diary Results
Teleoscoping differs from keyword searches. Participants found that the ranking system differed from a normal keyword search. For many participants, it took some time to (1) differentiate results of a keyword search from results of a Rank operation; and (2) differentiate valid results that did not meet their expectations due to the documents that existed in the corpus from invalid results due to bugs or mental model inconsistencies. There were negative transfer effects from being used to keyword searches which took multiple sessions to unlearn.
For example, P1 searched for “palliative” and was “…surprised by how many posts were about animals at end of life, which does not fit our defined research topic.” P1 then wished “…there was a feature that would take everything I had already put within one group and give me ‘more like this’”, which was exactly what the Rank operation was designed to do. Multiple participants corroborated this sentiment in their diaries (P5, P7, P8). This indicated a problem in the participants’ mental model, likely due to (1) how we were representing the Rank operation on the workspace; (2) our training, documentation and support materials; and (3) not enough time to learn the tool.
@@
Today I also spent time trying to go down the rabbit hole of different searches to try to gain a true appreciation for how this machine learning approach to data collection differs from just keyword searching within the Reddit search. This was really evident to me when I found a post where the OP had a palliative/terminal illness, and I wanted to find others where this was the case. I made a new folder for this category, then used the [Rank] feature, and immediately found one other post where the OP has cancer and was asking a friend to not mourn their death. It would be extremely difficult to keyword search for this type of topic, but it’s a very interesting and important area to capture (OPs with terminal illnesses). This was a great exploration! (P1)
This indicated that it was possible to learn the difference between keyword search and Rank, but that the learning curve was steep enough to require multiple hours of usage and documentation review.
@@
They further suggested adding a document quick viewer to aid in skimming. They then organized documents into groups, relabelling and changing the groups as they developed their understanding of the corpus and the tool. Then, they switched between using the Rank operation and the group feature to find relevant documents:
@@
After adding about half a dozed documents to the group ‘gay stuff’, I noticed that many of the documents are about gay panic. That is, the fear of being (wrongly or correctly) as gay, the dislike of anything perceived as gay, and an aversion from being around gay people. I changed the group name to ‘gay panic‘ to reflect this…once I had about 13–14 results, I opened the [Rank] window for the group. Looking at the first two pages of results, none of them that weren’t already in the group seemed very relevant, mostly judging by their titles and occasionally by the first couple of sentences in the document. I refined the search…[with a] couple of documents that I thought particularly demonstrated gay panic. (P7)
@@ 4.2.5. System Log Results
During the user study, the system maintained a log of user actions as they interacted with the system. Across the 5 participants, 7 sessions with the Rank operations were tracked (one participant created 3 separate sessions for themselves). The mean number of actions tracked per session was approximately 310 (median = 286, minimum = 97, maximum = 684). Actions included such things as session initialization, creation/movement/deletion of windows, keyword searches, instantiation of ranking and results, in short, any conditions where the state of the user workspace was altered.
The actions were logged and then visualized to better understand user interactions. Generally, users made use of an iterative process to find documents of interest, alternating between putting documents into groups and instantiating new Rank operations to find new documents relevant to their queries and then sorting them into their groups. The number of Rank operations created across the study were relatively small however (between 2 and 4).
@@ 4.2.6. Focus Group Results
Mental model need: Rank state needs to be inspectable. In our tested design, Rank state was coupled directly to a single group. When the group was updated, the Rank changed. Individual documents within the group could be selected to weight the Rank search vector closer to that document. This confused participants about the state of the Rank. We decided that the Rank inputs and outputs needed a more explicit visual representation and decoupled from a single group, allowing for multiple input sources.
@@
Mental model need: Clarify the exploration metaphors onscreen. For example, participants wondered whether they were being “dropped off in the landscape” of documents and “going down different paths” (P7). If so, they wanted a record of the paths and some way to compare paths directly onscreen. Participants agreed that “seeing it” and “understanding how close documents are to each other in space [is important].” (P1)
@@
Conceptual need: Confidence in path saturation. Participants agreed that they did not need to show total path exhaustion, rather, they needed a sense of saturation and demonstrable information power. Their imagination of the use case for Teleoscope was for data curation, which did not mean finding every piece of relevant data but instead finding enough representative samples of data.
Conceptual need: Explaining the methodology to reviewers. Participants were concerned about how to explain the Teleoscope process to reviewers at a high level, but were not concerned with the details of the Teleoscope process. As long as they had a clear metric that they could point reviewers to (e.g., a paper that describes the metric), they did not particularly care about which metric to use. This was a defensive publication strategy: in some qualitative papers, keyword searches are reported on directly to demonstrate that they had reached saturation or sufficient information power. Large data sets and less of a value of statistical rigour for reviewers means that any reasonable metric could be argued for and used.
Conceptual need: differentiating data curation from analysis. Participants were unsure whether using Teleoscope constituted a violation of rigour such that it mixed data curation and analysis research phases. Particularly if text-level analysis was to be supported, they felt that it might not be appropriate to allow for both in the interface at once. We found that participants used Teleoscope as we intended: they explored data in a manner that extended beyond keywords searches. For example, participants reported the following:
@@ 4.2.7. Post-Hoc Analysis and Recommendations from Visualization Group
4.2.8. Case Study 2: Conclusion
@@
  • CS2.RQ1 Result. Participants understood the overall concept of the Ranking operation and workspace as a whole, likening it to walking down different paths in an unknown region. Even though this metaphor came from the participants, the operation inputs and system state were unclear; therefore we decided that the path metaphor should be visualized as directly as possible. Participants did not use any collaboration features of Teleoscope, which may be due to a lack of impetus, our study environment, or because of our design. In the next study, we addressed these possibilities to ensure collaboration features were used.
  • CS2.RQ2 Result. Participants interpreted Teleoscope primarily as a data curation tool. Further, they were unsure about whether it infringes on analysis and had concerns about academic rigour if it does.
Figure 12. An example of a workflow from Case Study 3 our long-term deployment. Pictures are actual screenshots of research artifacts from our participant research team’s data-gathering phase. Participants worked both individually and collaboratively on the Teleoscope interface, and collaboratively on Google Docs, Zoom, and in-person. Due to the existing qualitative research culture in their research group, keyword searches were the focus of their data curation approach. ML features were used to discover new keywords, find thematically similar documents that did not have specific keywords, and to saturate groups with relevant documents.
4.3. Case Study 3: Field Deployment
After completing a redesign after Case Study 2, we were interested in a long-term focused field deployment with qualitative research groups who could bring their own research needs to us. We also wanted to move Teleoscope out of a simulated environment and into a production environment where real-world motivations and difficulties would be encountered.
Case Study 2 had many study contrivances, such as a research topic that none of the participants were using for their own research. We felt that it was important to see how Teleoscope would perform when participants were subjected to all the benefits, consequences and costs of real research. We also committed to ongoing feature development (and bug fixes) using a quasi co-design methodology.
@@ 4.3. Case Study 3: Field Deployment
  • CS3.RQ1. When put in a real-world environment, how did researchers incorporate Teleoscope into their own research practice? Which features, processes, and workflows within Teleoscope were commonly used and in what way?
  • CS3.RQ2. Using Teleoscope, to what extent were researchers able to feel confident that they were able to retrieve data using criteria that are important to qualitative researchers, i.e., richness, information power, saturation, etc.?
@@ 4.3.1. Participants
Three PIs and research groups were recruited; only one research group was able to commit to long-term use of Teleoscope. The final team was comprised the PI from Case Study 2 and a research team of two graduate RAs that were recruited specifically to use Teleoscope. Participants were not reimbursed for their time by us since they were being supported extensively by our design team for their own research project; our understanding is that RAs were compensated as normal by the PI.
@@ 4.3.3. Results
The research team used Teleoscope for data curation for their research project on nurses and structural inequality as articulated in Reddit’s r/nursing forum, which is where working nurses post about their day-to-day problems. Of their own initiative and in alignment with their own qualitative methodological approach, the team’s main conceptualization of data curation was via an external google document that contained a set of keywords to be used in searches (see Figure 12). After exploring the data using a variety of features in the Teleoscope workspace (described below), they would update their keyword list with successful and unsuccessful searches. Here are the ways in which the research team used the Teleoscope workspace features:
@@
Teleoscope helped with search saturation. The researchers reported (1) making groups from documents from their keyword searches; (2) piping the groups into the Ranking operation as controls; (3) determining which documents they had not yet read and (4) adding those documents to the groups. This helped to see the parts of the document space that they had not yet captured with a keyword search.
@@
Putting each one of our single groups like indigenous, vulnerable disabilities, [an RA] put them into the [Rank] and then basically went through to see like which ones we hadn’t read…We were just trying to expand our data set and be exhaustive. (P1)
@@
It was actually really helpful to start with keyword searches. And to be able to build out this groups structure, and then [Rank] from there. Whereas maybe if we’d had a more like drilled in topic, we could have just gone from there.(P1)
@@
Projections have potential to allay methodological concerns. In the focus group, it was brought up as a concern that using Teleoscope might be too close to analysis and confuse methodological rigour (in rigorous qualitative research, data curation and analysis stages are kept distinct so as not to predetermine results). However, the projection operation, which was added after the focus group, seemed to have potential to populate the interface with unexpected results:
@@
The speed and ease of use of the Projetion operation helped with researchers’ sense of completion as well:
@@
4.3.4. Case Study 3: Conclusion
@@
  • CS3.RQ1 Result. Researchers incorporated Teleoscope into their research practice as a data curation tool, working between an external Google doc for keywords and the interface itself. Teleoscope was used to explore parts of the document space where keywords would not be easy or obvious to find. The most common features used were the keyword search, document reading and grouping. [Ranks] were used after grouping. Projections were used after Teleoscoping and grouping.
@@
  • CS3.RQ2 Result. Teleoscope helped to provide confidence that a corpus was being more fully and rigorously explored by providing both ranked and randomized example documents.
5. Related Work
In this section, we discuss the similarity and differences between Teleoscope and other related works. With Teleoscope, we departed from standard topic modelling because we are not interested in creating categorical topics, but instead are interested in supporting researchers in the first stage of developing rich, specific themes. The difference between categories and themes is subtle, but it is important to qualitative researchers. Rather than creating deductive categories, qualitative researchers are interested in inductive interpretation.
The difference is illustrated well with Cody, Rietz et al.’s excellent qualitative coding support tool (Rietz and Maedche, 2021). Cody addresses the problem of annotating specific sentences with qualitative codes (similar to tagging) by connecting keywords with set operations to create inclusion/exclusion rules. These rules are then applied across a corpus. This supports a style of coding that is fundamentally categorical, that is, the main action of analysis is to organize the different sentences into code categories. This is helpful in inductive coding, and is seen in new LLM-supported tools for qualitative analysis and sense-making (Gao et al., 2024; Kim et al., 2024; Liu et al., 2023). Telescope, by contrast, is not attempting to support users in coding and labeling documents; instead we focus on curating document sets by visualizing provenance relationships.
On the categorical side, many topic modelling visualizations take the approach of acting directly on topic as document categories. For example, Serendip (Alexander et al., 2014) and TopicSifter (Kim et al., 2019) both approach the problem of topic exploration and discovery by allowing users to filter and/or drill down into topics. For these tools, the topics are the primary objects of interaction, and the goal is to explore the topics themselves. Both are a top-down approach. Teleoscope is most conceptually similar to Scholastic by Hong et al. (Hong et al., 2022) because we also take a human-centred approach and provide a hierarchical clustering algorithm as a “machine-in-the-loop” approach. Their elegant approach of visualizing the document space differs greatly from our choice of visualizing the process of discovery. Also, as with Cody, they focus on document coding, which we do not.
Teleoscope differs from topic modelling tools by focusing on (1) discovering relevant documents rather than making claims about a whole corpus; (2) supporting collaboration, reproducibility, and schema expression; (3) co-developing themes through creative data exploration; and (4) allowing users to interact in real time with a larger corpus than other tools (DG1, DG2, DG3).
6. Design Choices and Limitations